OpenStudy (icalibear):

Find the length of the midsegment? (I already have the answer, I need it checked)

OpenStudy (icalibear):



OpenStudy (anonymous):

is it a right triangle?

OpenStudy (icalibear):

It looks like it but my question doesnt say anything about that

OpenStudy (icalibear):

there's no right angle sign so no

OpenStudy (anonymous):

It doesn't need to be a right triangle. It's a question on proportions dealing with similar triangles. 6x + 4 = (1/2)(4x + 56) So, first, solve for "x".

OpenStudy (anonymous):

This equality comes from the 2 horizontal lines being parallel and bisecting the 2 sides of a triangle. Notice how they do so halfway down.

OpenStudy (anonymous):

Once you solve for "x", you can get your midsegment by just plugging that "x" back into 6x + 4

OpenStudy (icalibear):

I started solving the equation and ended up at the end with 10x = 26 which isnt right

OpenStudy (icalibear):

how do i find x im failing at it over and over

OpenStudy (anonymous):

6x + 4 = (1/2)(4x + 56) -> 6x + 4 = 2x + 28 4x = 24 -> x = 6 So, now, substitute back in.

OpenStudy (icalibear):

= 16

OpenStudy (icalibear):

thats not one of my options

OpenStudy (anonymous):

icalibear, you're just punching in numbers arbitrarily. Show us your work as you answer like tcarroll did and then we can tell you where or if you made a mistake.

OpenStudy (icalibear):

6x + 4 = (1/2)(4x + 56) 6x + 4 = 2x + 28 4x + 4 = 28 4x = 24 x = 6 6(6) + 4 = 16

OpenStudy (anonymous):

I already gave you: x = 6 All you have to do is: 6(6) + 4 for your answer. You have almost no work left to do.

OpenStudy (icalibear):

I got 16 as shown above. But that's not one of my options. So what do I do next?

OpenStudy (anonymous):

What is (6)(6) ?

OpenStudy (icalibear):

oh 16.i was thinking 12. i dont know why,

OpenStudy (icalibear):

36*

OpenStudy (icalibear):

not 16

OpenStudy (anonymous):

Can you calculate (6)(6) + 4 for your correct answer?

OpenStudy (icalibear):

its 40 i was just adding instead of multiplying for some reason

OpenStudy (icalibear):

thats where i went wrong lol thanks for helping :P

OpenStudy (anonymous):

np You're welcome